Find the consecutive odd integers such that the sum of the least integer and the middle integer is 49 more than the greatest integer.

Answer by ikleyn(53751)

.

Let x be the smallest odd number of these three.


Then the integers are x, (x+2) and (x+4).


The condition requires


    x + (x+2) = (x+4) + 49.


Simplify and find x.


    2x + 2 = x + 53

    2x - x = 53 - 2

    x      = 51.


ANSWER.  The three numbers are  51, 53 and 55.

Solved.